# STRelay: A Universal Spatio-Temporal Relaying Framework for Location Prediction with Future Spatiotemporal Contexts

### v1

@@ Conclusion
We developed our model using the PyTorch framework and conducted experiments on the following hardware platform (CPU: Intel(R) Xeon(R) Gold 5320, GPU: NVIDIA GeForce RTX 3090). We evaluate STRealy and the baselines in the location prediction task. We empirically set the dimension of hidden states and all (POI, timestamp, and user) embedding sizes as 10, and the temporal decay factor $\alpha=0.1$ and spatial decay factor $\beta=100$ on all datasets, and the sequence length is 20. All baseline implementations are either provided by the original authors or based on the original papers.
Additional Results on Overall Performance Comparison
@@
Table 4 shows the results of evaluation metrics Acc@5 and NDCG@5 overall location prediction performance comparison. We observe that STRelay+Graph-Flashback consistently outperforms all baselines, achieving 6.63% and 5.84% improvement on average over the best-performing baselines across four datasets on Acc@5 and NDCG@5, respectively. Additionally, our STRelay framework consistently enhances the performance of corresponding base prediction methods STGN, Flashback, SNPM, and Graph-Flashback with an average improvement of 2.54%-10.9% across four datasets.
Additional Results on Ablation Study
Table 5 shows the ablation experiment results of evaluation metrics on Acc@5 and NDCG@5. First, the complete STRealy framework outperforms the spatial-ablated variant (w/o spatial) and temporal-ablated variant (w/o temporal) by 2.96% and 2.52% across four base models and datasets in most cases. Second, we observe that the complete STRelay framework outperforms the w/o relaying variant by 1.57%-2.86% performance improvement across four base models and datasets.
@@
Additional Results on Impact of Spatiotemporal Context Granularity
@@
Firstly, to investigate the impact of temporal context granularity, we train the model with future time intervals ranging from 0.5 to 24 hours. We report the results of Acc@10 on four datasets as shown in Figure 6. We observe a general trend where performance decreases as the temporal granularity becomes coarser. Besides, we find that the performance of the 1h granularity achieves the best results across four datasets. Based on the above observations, we adopt 1h as the optimal temporal granularity.
Secondly, to investigate the impact of spatial context granularity, we train the model with spatial intervals ranging from 0.5 km to 30 km. The corresponding results of Acc@10 are presented in Figure 7. The results demonstrate the same overall trend observed for temporal granularity, where performance decreases as spatial granularity becomes coarser. Moreover, while the 1km spatial granularity yields slightly lower performance than 0.5km in the Singapore and Tokyo datasets, it achieves optimal performance in the Istanbul and Moscow datasets; thus, we adopt 1km as our spatial granularity in our experiment.

### v2

@@ 5.1.4 Experiment Settings
We developed our model using the PyTorch framework and conducted experiments on the following hardware platform (CPU: Intel(R) Xeon(R) Gold 5320, GPU: NVIDIA GeForce RTX 3090). We evaluate STRealy and the baselines in the location prediction task. The implementations of all baselines and base models are either provided by the original authors or based on the original papers.
5.2 Overall Performance Comparison
Table II and Table III show the overall performance on all four datasets, and for each dataset and each metric, boldface indicates the best overall result, while underline highlights the strongest baseline method. We observe that among all STRelay enhanced methods, the STRelay+Graph-Flashback achieves the best performance in most cases, except on NDCG and MRR on the Tokyo dataset, where the STRelay+LoTNext achieves slightly better performance than STRelay+Graph-Flashback. More importantly, compared with the best-performing baselines, STRelay+Graph-Flashback yields average relative improvements of 6.72%, 5.92%, 3.74%, and 4.45% on the Istanbul, Tokyo, Singapore, and Moscow datasets, respectively, demonstrating its effectiveness. Moreover, DeepMove, STAN and GETNext consider future temporal information but still yield suboptimal performance compared to STRelay enhanced baselines.
@@ 5.2 Overall Performance Comparison
Notably, we observe that our STRelay framework consistently improves the location prediction performance of the corresponding base prediction methods STGN, Flashback, Graph-Flashback, SNPM, and LoTNext, on all datasets, which demonstrates the universal effectiveness of the STRelay framework. STRelay boosts the performance of the above five baselines with an average improvement of 2.49%-11.30% across different datasets in the next location prediction task.
In the following, we conduct a systematic ablation study below to verify the usefulness of different components of our STRelay framework.
5.3 Ablation Study
We consider four variants to quantify their impact on location prediction performance and demonstrate their utility.
  • w/o spatial is a variant that removes future spatial context. It is also equivalent to only considering future temporal contexts for location prediction.
  • w/o temporal is a variant that removes future temporal context. It is also equivalent to only considering future spatial contexts for location prediction.
  • w/o relaying is a variant that models future temporal and spatial contexts not in a relaying manner, but in a parallel manner. Specifically, we decoupled spatial and temporal context modeling (i.e., removing the sequential dependency where spatial modeling conditions on temporal modeling).
  • w/o spatiotemporal is a variant that removes both future spatial and temporal contexts, reducing STRelay to the plain base model.
The ablation results are reported in Tables IV and V. First, we observe that the complete STRelay framework consistently outperforms the spatial-ablated variant (w/o spatial) by 2.65% on average across five base models and four datasets, highlighting the essential contributions of modeling future spatial context. Meanwhile, we observe that the w/o temporal variant (which retains only future spatial contexts) outperforms the w/o spatiotemporal variant by 3.89% on average across all five base models and four datasets, further validating the usefulness of modeling future spatial context.
Similarly, compared to the complete STRelay framework, the temporal-ablated variant (w/o temporal) that removes future temporal contexts, leads to an average performance drop of 2.31%, confirming the critical importance of future temporal context. The usefulness of modeling future temporal context is also validated by the w/o spatial variant (which retains only future temporal context) outperforms the w/o spatiotemporal variant by 3.46% on average across all five base models and four datasets.
Moreover, we notice that the complete STRelay framework outperforms the w/o relaying variant by 0.47%-2.90% performance improvement across five base models and datasets, which demonstrates that our learning in a relaying manner indeed boosts performance by flexibly capturing the inherent relationship between temporal and spatial contexts.
Finally, the complete STRelay framework outperforms the context-free variant (w/o spatiotemporal) by 6.28% on average across all five base models and four datasets. This validates our core motivation, i.e., future spatiotemporal contexts provide highly informative signals for next-location prediction, and demonstrates the superiority of the proposed STRelay framework.
Figure 4: Performance improvement on different location types
Figure 5: Performance improvement on different user groups
5.4 Performance Improvement Analysis on Location Types and User Groups
First, to study the performance improvement across different location types, we classify locations into corresponding activity categories following previous work [33]. Specifically, we adopt a binary classification scheme consisting of: leisure-related locations (e.g., Outdoors and Entertainment) and daily-routine-related locations (e.g., College and Residence). Figure 4 shows the results of the percentage of performance improvement of Acc@10 in predicting these two types of locations on the four datasets. We observe that the performance improvements stemming from leisure-related locations are consistently and significantly higher than those from daily-routine-related locations by 2.66%-4.57%, which demonstrates that modeling future spatiotemporal contexts is particularly beneficial for entertainment-related activities, which are more uncertain than daily-routine-related activities.
Second, we categorize users into two distinct groups based on their radius of gyration [62, 59], which is calculated as the root mean squared distance of all locations in a trajectory from the geographical center of the trajectory, and obtain two groups: long-distance preferred users (top 50%) and short-distance preferred users (bottom 50%). We report the performance improvements on these two user groups. As shown in Figure 5, both user groups exhibit performance improvements in Acc@10 across all four datasets, with long-distance preferred users consistently demonstrating larger improvement (0.25%-1.07% higher than short-distance preferred users). These results indicate that modeling future spatiotemporal contexts is more effective in capturing the mobility patterns of users who prefer traveling longer distances, which often implies higher uncertainty for next locations.
@@ 5.4 Performance Improvement Analysis on Location Types and User Groups
5.5 Impact of Spatiotemporal Context Granularity
@@
Firstly, to investigate the impact of temporal context granularity, we train the model with future time intervals ranging from 0.5 to 24 hours. We report the results of Acc@10 and NDCG@10 on four datasets as shown in Figure 6. We observe that both metrics exhibit a clear declining trend as the temporal granularity becomes coarser (i.e., longer intervals lead to poorer performance). Notably, the 1-hour granularity consistently yields the highest Acc@10 and NDCG@10 on all datasets. Based on these observations, we adopt 1 hour as the optimal temporal granularity in our experiment settings.
Secondly, to investigate the impact of spatial context granularity, we train the model with spatial intervals ranging from 0.5 km to 30 km. The corresponding results of Acc@10 and NDCG@10 are presented in Figure 7. Both metrics follow the same trend as observed for temporal granularity: the performance degrades as the spatial interval becomes coarser. Although the 0.5 km setting performs slightly better on the Singapore and Tokyo datasets, the 1 km granularity delivers the highest scores on the Istanbul and Moscow datasets, while achieving the best NDCG@10 across all four datasets. Therefore, we select 1 km as the optimal spatial granularity in our experiment settings.
6 Conclusion
In this paper, we propose STRelay, a universal framework designed to explicitly model the future spatiotemporal contexts in a relaying manner; it can be flexibly integrated with different location prediction models to boost the prediction performance. Experimental results show that STRelay consistently boosts the performance of state-of-the-art baselines with an average improvement of 2.49%-11.30% across four datasets, demonstrating the universal effectiveness of STRelay. Moreover, we find that the future spatiotemporal contexts are particularly helpful for entertainment-related locations and also for user groups who prefer traveling longer distances. The performance gain on such non-daily-routine activities, which often suffer from higher uncertainty, is indeed complementary to the base location prediction models that excel at modeling regular daily routine patterns.